# Comprehensive X-ray Observations of the Exceptional Ultra-long X-ray and Gamma-ray Transient GRB 250702B with Swift, NuSTAR, and Chandra: Insights from the X-ray Afterglow Properties

### v2

@@ 4.2 X-ray Luminosity and Prompt-Afterglow Correlations
Figure 11: Rest frame X-ray luminosity (0.30.3-$10$ keV) at 11 hours versus gamma-ray energy (11-$10,000$ keV). GRB 250702B is shown by a red star. We measure 11 hours from the start of the “D” burst. A sample of short (black squares) and long GRBs (gray circles) are compiled from Nysewander2009; E. Berger (2014). We show GRB 221009A as a purple square (OConnor2023; Lesage2023; E. Burns et al., 2023). A sample of ultra-long GRBs is shown for comparison (GRBs 101225A; 111209A, 121027A, 130925A; Levan2014; P. A. Evans et al., 2014a). The long GRB correlation of $L_{\textrm{X,11}}\approx$$5.6\times 10^{44}(E_{\gamma,\textrm{iso}}/10^{51}\,\textrm{erg})^{0.72}$ erg s-1 is shown by a black dashed line (E. Berger, 2014). Reproduced from OConnor2025.
@@
The agreement of GRB 250702B with this correlation depends on the exact onset time (Figure 1), as adopting an earlier start time (e.g., the initial EP detection) is likely to modify the X-ray luminosity. Unfortunately, we do not have information on the X-ray lightcurve of GRB 250702B during its prompt gamma-ray emitting phase. In Figure 11, we have assumed the “D” burst as $T_{0}$. Utilizing the “E” burst as $T_{0}$ does not modify this conclusion. The strong agreement with the $E_{\gamma,\textrm{iso}}$-$L_{\textrm{X,11}}$ correlation circumstantially favors these start times over the initial EP detection, as in Figure 1.
@@ 4.3 An Internal or External Origin for the X-ray Lightcurve?
In the second scenario for the X-ray emission, the external X-ray emission does not lie very far below the accretion driven excess that produces the flaring behavior (Figure 9). Our afterglow modeling shows that this scenario is easy to produce with reasonable kinetic energies (Figure 20) with the forward shock emission beginning to dominate the observed lightcurve at >>22 days. It is important to note that the X-ray variability observed for GRB 250702B (Figure 4) is at a lower amplitude, and much less frequent, than observed in known relativistic TDEs (e.g., Sw J1644+57; D. N. Burrows et al. 2011; J. S. Bloom et al. 2011). The current sample of relativistic TDEs shows dramatic ($10\times$) dips in the flux on short timescales, which are not observed here. There is minor evidence for a dip at the end of the initial NuSTAR observation (Figure 4) and at $\sim$$2.2$ d (observer frame) by Swift (Figure 1), but the variability amplitude is much smaller than in Sw J1644+57 or AT2022cmc (I. Andreoni et al., 2022; Pasham2023). This could be due to the proximity of the external afterglow to the flux level of the accretion driven emission at $\sim$22 days (observer frame).
@@ 4.4 Results of External Afterglow Modeling
The long-lasting prompt phase adds difficulty in determining the precise onset time of the afterglow. Corroborated by our temporal analysis of the X-ray lightcurve (§3.1.1; Figure 1), we performed our afterglow modeling relative to the GBM “E” burst (MJD 60858.682; Table 2). We support interpreting this as the onset time by invoking a picture where each independent GBM trigger (or prompt episode) is due to an internal shell-shell collision and essentially marks the launch time of an ultra-relativistic shell. As each shell collides, they merge to form a larger, more energetic shell, which eventually catches up to, and then refreshes, the blast wave producing the forward shock (SariMeszaros2000; KumarPanaitescu2000). This refreshed shock then resets the inferred afterglow $T_{0}$ with the most energetic shell (the “E” burst) dominating the afterglow. The results of our afterlgow modeling using the “E” burst as $T_{0}$ are shown in Figures 9 and 20. However, we note that utilizing the “D” burst as $T_{0}$ only marginally changes our results for the afterglow parameters, see Figures 22 and 21.
@@ 4.5 Implications from the Gas-to-Dust Ratio
Kruhler2011dust interprets the lower $N_{\textrm{H,z}}/A_{\textrm{V,z}}$ ratio as two independent absorbers with local ionized metals in the circumburst environment producing the inferred X-ray absorption $N_{\textrm{H,z}}$ and a physically separate dusty absorber in the host galaxy producing the high visual extinction $A_{\textrm{V,z}}$. This scenario is plausible for GRB 250702B and is consistent with the independent inference of a dusty galaxy based on the host spectral energy distribution modeling of J. Carney et al. (2025). However, we cannot rule out that the local environment of GRB 250702B simply differs from classical GRBs and has SMC-like gas-to-dust properties (P. Bouchet et al., 1985). We note that there are existing selection effects against the discovery of the most highest extinguished GRBs, and that rapid, deep near-infrared imaging is required to identify other potential members of this class.
@@ 5.1 An Intermediate-mass Black Hole Progenitor
We therefore see that a white dwarf disruption would produce a timescale too short to be consistent with the observed, day-long duration associated with the prompt emission. Specifically, noting that the fallback rate peaks on a timescale of $\sim 0.16$ hr for a $4\times 10^{3}M_{\odot}$ black hole and using the $\propto M_{\bullet}^{1/2}$ scaling implied by Equation (3) (and shown in the right panel of Figure 13), a black hole mass of $M_{\bullet}\simeq 4\times 10^{7}M_{\odot}$ would be required to yield a peak timescale of $\sim$$\big(\frac{1}{1+z}\big)1.5$ days. This is well in excess of the Hills mass for the disruption of a white dwarf, which is comparable to $\sim a\,few\times 10^{5}M_{\odot}$ if the black hole is spinning and the orbital angular momentum of the star is aligned with the spin axis of the black hole. While the fallback rates shown in Figure 13 (right panel) are obtained from the complete disruption of a white dwarf, a partial disruption could yield a longer peak timescale by a factor of at most a few [^14] Figure 2 of D. Garain & T. Sarkar (2024) shows fallback rates for the disruption of a $1M_{\odot}$ white dwarf by a $10^{3}M_{\odot}$ black hole for a range of orbital pericenter distances, wherein the peak timescale shows very little variation over the entire range.. A longer peak timescale could also arise from the white dwarf being on an unbound (eccentricity >1>1) orbit about the black hole, but, as shown in Figure 4 of M. Cufari et al. (2022), $t_{\rm peak}$ changes by less than a factor of $\sim 10$ over a wide range of orbital eccentricities, implying that the black hole mass required to generate a fallback rate having a peak timescale of at least a day would still exceed the Hills mass for the disruption of a white dwarf. A white dwarf does not produce a long lasting enough central engine as required by the prompt phase unless a tightly bound orbit (in which essentially all debris is accreted on the orbital timescale) combined with (rare) repeated partial disruptions are considered (Levan2025). Such scenarios likely require a (likely rare) Hills capture mechanism (J. G. Hills, 1988). Based on these arguments we strongly disfavor the disruption of a white dwarf by an intermediate-mass black hole as the progenitor of GRB 250702B.
@@ 5.2 A Stellar-mass Black Hole Progenitor
We additionally note that while the exact progenitor scenario is uncertain, the existence of a supernova would be a critical “smoking gun” for a stellar explosion and therefore a stellar-mass black hole engine. A supernova is expected in both a classical and blue supergiant collapsar scenario (e.g., Nakauchi2013) as well as in a helium merger (C. L. Fryer & S. E. Woosley, 1998; Neights2025), but not in relativistic TDE scenario (§5.1). Future observations with the James Webb Space Telescope will be critical to provide robust constraints on this scenario (see also B. P. Gompertz et al., 2025).
@@ 6 Conclusions
We demonstrate that the overall characteristics of GRB 250702B do not cleanly match either known relativistic TDEs or ultra-long GRBs. With the exception of the initial soft X-ray detection a full day before the initial gamma-ray triggers, the properties have a closer match to those seen by ultra-long GRBs, but require an even longer lived central engine ($\gtrsim$33 days). In particular, the early soft X-ray detections (H. Q. Cheng et al., 2025a; EP250702a-arxiv) are not easily explained in the standard collapsar model (Woosley1993). As such, we favor a hybrid stellar-mass black hole progenitor, such as a micro-TDE (P. Beniamini et al., 2025), though see also Neights2025 and J. Granot et al. (2025) for alternative explanations.
The “smoking gun” evidence to distinguish between the main progenitor scenarios is currently missing, but may be obtained through sensitive late-time X-ray observations. As the exact nature of the source is debated, continued X-ray monitoring of the event to monitor for a possible jet shutoff is critical to fully constrain the nature of the source. Deep spectroscopic observations of the host galaxy environment can likely provide additional cues (e.g., the local environment’s metallicity and star formation rate) to distinguish between the possible progenitor scenarios. Additional observations of similar transients, hopefully discoverable via the Einstein Probe (Yuan2025), will aid in uncovering the true nature of GRB 250702B.
@@
The authors acknowledge Karl Forster and the NuSTAR SOC for their rapid scheduling of our observations, as well as Pat Slane and the CXO staff for approval and scheduling of our Chandra DDT request. B. O. acknowledges Brian Grefenstette, Oli Roberts, Eliza Neights, and Eric Burns for useful discussions. B. O. thanks Ben Gompertz, Andrew Levan, and Antonio Martin-Carrillo for sharing the redshift derived by JWST. B. O. is supported by the McWilliams Postdoctoral Fellowship in the McWilliams Center for Cosmology and Astrophysics at Carnegie Mellon University. J. D. acknowledges support from NASA under award numbers NAS5-0136, 80NSSC25K7567, and 80NSSC25K7788. D. R. P acknowledges support from NASA in the form of a NuSTAR guest observer program. E. R. C. acknowledges support from NASA through the Astrophysics Theory Program, grant 80NSSC24K0897. A. B. acknowledges support from NASA through the FINESST program, grant 80NSSC24K1548. M. M. is supported by an appointment to the NASA Postdoctoral Program at the NASA Goddard Space Flight Center, administered by Oak Ridge Associated Universities under contract with NASA. J. H. acknowledges support from NASA under award number 80GSFC21M0002. P. B. work was funded by a grant (no. 2020747) from the United States-Israel Binational Science Foundation (BSF), Jerusalem, Israel and by a grant (no. 1649/23) from the Israel Science Foundation. M. B. is supported by a Student Grant from the Wübben Stiftung Wissenschaft. Funded in part by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2094 – 390783311. This work used resources on the Vera Cluster at the Pittsburgh Supercomputing Center (PSC). Vera is a dedicated cluster for the McWilliams Center for Cosmology and Astrophysics at Carnegie Mellon University. We thank the PSC staff for their support of the Vera Cluster. This work made use of data from the NuSTAR mission, a project led by the California Institute of Technology, managed by the Jet Propulsion Laboratory, and funded by the National Aeronautics and Space Administration. This research has made use of the NuSTAR Data Analysis Software (NuSTARDAS) jointly developed by the ASI Space Science Data Center (SSDC, Italy) and the California Institute of Technology (Caltech, USA). The scientific results reported in this article are based on observations made by the Chandra X-ray Observatory. This research has made use of software provided by the Chandra X-ray Center (CXC) in the application package CIAO. This work made use of data supplied by the UK Swift Science Data Centre at the University of Leicester. This research has made use of the XRT Data Analysis Software (XRTDAS) developed under the responsibility of the ASI Science Data Center (ASDC), Italy. This research has made use of data and/or software provided by the High Energy Astrophysics Science Archive Research Center (HEASARC), which is a service of the Astrophysics Science Division at NASA/GSFC.
@@ Appendix A Log of Observations
Table 1: Log of X-ray observations of EP250702a obtained with NuSTAR (PI: Pasham) and Chandra (PI: O’Connor, ObsID: 31011; and PI: Eyles-Ferris, ObsID: 31468). The fluxes are reported in 33-$79$ keV for NuSTAR and 0.30.3-$10$ keV for Chandra. The days from discovery, using the trigger time of GRB 250702D, is provided in the $\delta T$ column.
@@ Appendix B Log of Trigger Times
Einstein Probe EP250702aaaEarlier detections on July 1 were obtained by EP (Hui Sun, private communication), but this is the earliest reported detection time (H. Q. Cheng et al., 2025a). 2025-07-02 02:53:44 60858.1206 H. Q. Cheng et al. (2025a)
@@ Appendix C Temporal Fit Results
Figure 14: Top Left: Results of our temporal fit with the onset time free. The fit includes all XRT and Chandra data. Top Right: Results of our temporal fit with the onset time fixed to GRB 250702D’s trigger time. Bottom Left: Same as top left, but only including XRT data after 0.55 days. Bottom Right: Same as bottom right, but only including XRT data after 0.55 days. The posterior distributions on onset time are also shown in Figure 1.
Figure 15: Results of our temporal fit with a broken powerlaw with the onset time free.

### v3

@@ 4.2 X-ray Luminosity and Prompt-Afterglow Correlations
Figure 11: Rest frame X-ray luminosity (0.30.3-$10$ keV) at 11 hours (in the rest frame) from discovery versus gamma-ray energy (11-$10,000$ keV). GRB 250702B is shown by a red star. We measure 11 hours from the start of the “D” burst (discovery trigger). We show Sw J1644+57 as a shaded region (see text for details; D. N. Burrows et al. 2011). A sample of short (black squares) and long GRBs (gray circles) are compiled from Nysewander2009; E. Berger (2014). We show GRB 221009A as a purple square (OConnor2023; Lesage2023; E. Burns et al., 2023). A sample of ultra-long GRBs is shown for comparison (GRBs 101225A; 111209A, 121027A, 130925A; Levan2014; P. A. Evans et al., 2014a). The long GRB correlation of $L_{\textrm{X,11}}\approx$$5.6\times 10^{44}(E_{\gamma,\textrm{iso}}/10^{51}\,\textrm{erg})^{0.72}$ erg s-1 is shown by a black dashed line (E. Berger, 2014). Reproduced from OConnor2025.
@@
In Figure 11, we consider the X-ray luminosity at 11 hours (rest frame) from the discovery of each transient. Thus, for GRB 250702B we have adopted the “D” burst as $T_{0}$. However, the agreement of GRB 250702B with this correlation depends on the exact afterglow onset time (Figure 1), as adopting an earlier start time (e.g., the initial EP detection) is likely to modify the X-ray luminosity (towards brighter values). In any case, we do find that utilizing the “E” burst as $T_{0}$ does not modify this conclusion.
@@ 4.3 An Internal or External Origin for the X-ray Lightcurve?
In the second scenario for the X-ray emission, the external X-ray emission does not lie very far below the accretion driven excess that produces the flaring behavior (Figure 9). Our afterglow modeling shows that this scenario is easy to produce with reasonable kinetic energies (Figure 20) with the forward shock emission beginning to dominate the observed lightcurve at >>22 days. It is important to note that the X-ray variability observed for GRB 250702B (Figure 4) is at a lower amplitude, and much less frequent, than observed in known relativistic TDEs (e.g., Sw J1644+57; D. N. Burrows et al. 2011; J. S. Bloom et al. 2011). The current sample of relativistic TDEs shows dramatic (up to $10\times$) dips in the flux on short timescales, which are not observed here. There is minor evidence for a dip at the end of the initial NuSTAR observation (Figure 4) and at $\sim$$2.2$ d (observer frame) by Swift (Figure 1), but the variability amplitude is significantly smaller than in Sw J1644+57 (J. S. Bloom et al., 2011; D. N. Burrows et al., 2011) or AT2022cmc (I. Andreoni et al., 2022; Pasham2023). However, we point out that the jetted TDEs Sw J2058+0516 (S. B. Cenko et al., 2012) and Sw J1112-8238 (G. C. Brown et al., 2015) do not show as pronounced dips as in Sw J1644+57 (J. S. Bloom et al., 2011; D. N. Burrows et al., 2011). The cause of this in GRB 250702B could be due to the proximity of the external afterglow to the flux level of the accretion driven emission at $\sim$22 days (observer frame).
@@ 4.4 Results of External Afterglow Modeling
The long-lasting prompt phase adds difficulty in determining the precise onset time of the afterglow. Corroborated by our temporal analysis of the X-ray lightcurve (§3.1.1; Figure 1), we performed our afterglow modeling relative to the GBM “E” burst (MJD 60858.682; Table 2). We support interpreting this as the afterglow onset time by invoking a picture where each independent GBM trigger (or prompt episode) is due to an internal shell-shell collision and essentially marks the launch time of an ultra-relativistic shell. As each shell collides, they merge to form a larger, more energetic shell, which eventually catches up to, and then refreshes, the blast wave producing the forward shock (SariMeszaros2000; KumarPanaitescu2000). This refreshed shock then resets the inferred afterglow $T_{0}$ with the most energetic shell (the “E” burst) dominating the afterglow. The results of our afterglow modeling using the “E” burst as $T_{0}$ are shown in Figures 9 and 20. However, we note that utilizing the “D” burst as $T_{0}$ only marginally changes our results for the afterglow parameters, see Figures 22 and 21.
@@ 4.5 Implications from the Gas-to-Dust Ratio
Kruhler2011dust interprets the lower $N_{\textrm{H,z}}/A_{\textrm{V,z}}$ ratio as two independent absorbers with local ionized metals in the circumburst environment producing the inferred X-ray absorption $N_{\textrm{H,z}}$ and a physically separate dusty absorber in the host galaxy producing the high visual extinction $A_{\textrm{V,z}}$. This scenario is plausible for GRB 250702B and is consistent with the independent inference of a dusty galaxy based on the host spectral energy distribution modeling of J. Carney et al. (2025). However, we cannot rule out that the local environment of GRB 250702B simply differs from classical GRBs and has SMC-like gas-to-dust properties (P. Bouchet et al., 1985). We do caution that $N_{\textrm{H,z}}/A_{\textrm{V,z}}$ ratios (e.g., T. Güver & F. Özel, 2009) are calibrated in regions where $A_{\textrm{V}}$<<11 mag and they may not apply in denser or dustier environments like observed for GRB 250702B. Regardless of the interpretation, we can conclude that all events with high $A_{\textrm{V,z}}$ fall in the same region of the parameter space, which is worth further investigation. We note that there are existing selection effects against the discovery of the most highest extinguished GRBs, and that rapid, deep near-infrared imaging is required to identify other potential members of this class, which will aid in understanding the origin of these small gas-to-dust ratios.
@@ 5.1 An Intermediate-mass Black Hole Progenitor
We note that the hydrodynamic simulation shown in Figure 13 (right panel) provides the fallback rate, which is not necessarily equal to the accretion rate, which is ultimately responsible for liberating the observed energy (by assumption here). In particular, there will be a substantial difference between the two if the viscous timescale is substantially longer than the timescale over which the material is supplied to the black hole, i.e., the fallback time (rees88). For standard $\alpha$-disc models and typical values of $\alpha$, the viscous time is generally short compared to the fallback time, and the timescale over which the viscous delays becomes significant is $\sim$ years (J. K. Cannizzo et al., 1990; lodato11). Additionally, recent works (e.g., mockler19; nicholl22) have found that the early-time luminosities of observed TDEs can be reproduced well with fallback models that impose zero viscous delay (though late-time observations of TDEs exhibit sustained levels of accretion that are consistent with a viscously spreading disc; vanvelzen19).
Based on Figure 13, we see that a white dwarf disruption would produce a timescale too short to be consistent with the observed, day-long duration associated with the prompt emission. Specifically, noting that the fallback rate peaks on a timescale of $\sim 0.16$ hr for a $4\times 10^{3}M_{\odot}$ black hole and using the $\propto M_{\bullet}^{1/2}$ scaling implied by Equation (3) (and shown in the right panel of Figure 13), a black hole mass of $M_{\bullet}\simeq 4\times 10^{7}M_{\odot}$ would be required to yield a peak timescale of $\sim$$\big(\frac{1}{1+z}\big)1.5$ days. This is well in excess of the Hills mass for the disruption of a white dwarf, which is comparable to $\sim a\,few\times 10^{5}M_{\odot}$ if the black hole is spinning and the orbital angular momentum of the star is aligned with the spin axis of the black hole. While the fallback rates shown in Figure 13 (right panel) are obtained from the complete disruption of a white dwarf, a partial disruption could yield a longer peak timescale by a factor of at most a few [^17] Figure 2 of D. Garain & T. Sarkar (2024) shows fallback rates for the disruption of a $1M_{\odot}$ white dwarf by a $10^{3}M_{\odot}$ black hole for a range of orbital pericenter distances, wherein the peak timescale shows very little variation over the entire range.. A longer peak timescale could also arise from the white dwarf being on an unbound (eccentricity >1>1) orbit about the black hole, but, as shown in Figure 4 of M. Cufari et al. (2022), $t_{\rm peak}$ changes by less than a factor of $\sim 10$ over a wide range of orbital eccentricities, implying that the black hole mass required to generate a fallback rate having a peak timescale of at least a day would still exceed the Hills mass for the disruption of a white dwarf. A white dwarf does not produce a long lasting enough central engine as required by the prompt phase unless a tightly bound orbit (in which essentially all debris is accreted on the orbital timescale) combined with (rare) repeated partial disruptions are considered (Levan2025). Such scenarios likely require a (likely rare) Hills capture mechanism (J. G. Hills, 1988). Based on these arguments we strongly disfavor the disruption of a white dwarf by an intermediate-mass black hole as the progenitor of GRB 250702B.
@@ 6 Conclusions
We demonstrate that the overall characteristics of GRB 250702B do not cleanly match either known relativistic TDEs or ultra-long GRBs. With the exception of the initial soft X-ray detection a full day before the initial gamma-ray triggers, the properties have a closer match to those seen by ultra-long GRBs, but require an even longer lived central engine ($\gtrsim$33 days). In particular, the early soft X-ray detections (H. Q. Cheng et al., 2025a; EP250702a-arxiv) are not easily explained in the standard collapsar model (Woosley1993). As such, we favor a hybrid stellar-mass black hole progenitor, such as a micro-TDE (P. Beniamini et al., 2025), though see also Neights2025 and J. Granot et al. (2025) for alternative explanations. Additional observations of similar transients, hopefully discoverable via the Einstein Probe (Yuan2025), will aid in uncovering the true nature of GRB 250702B.
@@
The authors thank the anonymous referee for their helpful comments that improved the clarity of the manuscript. The authors acknowledge Karl Forster and the NuSTAR SOC for their rapid scheduling of our observations, as well as Pat Slane and the CXO staff for approval and scheduling of our Chandra DDT request. B. O. acknowledges Brian Grefenstette, Oli Roberts, Eliza Neights, and Eric Burns for useful discussions. B. O. thanks Ben Gompertz, Andrew Levan, and Antonio Martin-Carrillo for sharing the redshift derived by JWST. B. O. is supported by the McWilliams Postdoctoral Fellowship in the McWilliams Center for Cosmology and Astrophysics at Carnegie Mellon University. J. D. acknowledges support from NASA under award numbers NAS5-0136, 80NSSC25K7567, and 80NSSC25K7788. D. R. P acknowledges support from NASA in the form of a NuSTAR guest observer program. E. R. C. acknowledges support from NASA through the Astrophysics Theory Program, grant 80NSSC24K0897. A. B. acknowledges support from NASA through the FINESST program, grant 80NSSC24K1548. M. M. is supported by an appointment to the NASA Postdoctoral Program at the NASA Goddard Space Flight Center, administered by Oak Ridge Associated Universities under contract with NASA. J. H. acknowledges support from NASA under award number 80GSFC21M0002. P. B. work was funded by a grant (no. 2020747) from the United States-Israel Binational Science Foundation (BSF), Jerusalem, Israel and by a grant (no. 1649/23) from the Israel Science Foundation. M. B. is supported by a Student Grant from the Wübben Stiftung Wissenschaft. Funded in part by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2094 – 390783311. This work used resources on the Vera Cluster at the Pittsburgh Supercomputing Center (PSC). Vera is a dedicated cluster for the McWilliams Center for Cosmology and Astrophysics at Carnegie Mellon University. We thank the PSC staff for their support of the Vera Cluster. This work made use of data from the NuSTAR mission, a project led by the California Institute of Technology, managed by the Jet Propulsion Laboratory, and funded by the National Aeronautics and Space Administration. This research has made use of the NuSTAR Data Analysis Software (NuSTARDAS) jointly developed by the ASI Space Science Data Center (SSDC, Italy) and the California Institute of Technology (Caltech, USA). The scientific results reported in this article are based on observations made by the Chandra X-ray Observatory. This research has made use of software provided by the Chandra X-ray Center (CXC) in the application package CIAO. This work made use of data supplied by the UK Swift Science Data Centre at the University of Leicester. This research has made use of the XRT Data Analysis Software (XRTDAS) developed under the responsibility of the ASI Science Data Center (ASDC), Italy. This research has made use of data and/or software provided by the High Energy Astrophysics Science Archive Research Center (HEASARC), which is a service of the Astrophysics Science Division at NASA/GSFC.
@@ Appendix A Log of Observations
Table 1: Log of X-ray observations of EP250702a obtained with NuSTAR (PI: Pasham) and Chandra (PI: Li, ObsID: 31003; PI: O’Connor, ObsID: 31011; PI: Eyles-Ferris, ObsID: 31468). The fluxes are reported in 33-$79$ keV for NuSTAR and 0.30.3-$10$ keV for Chandra. The days from discovery, using the trigger time of GRB 250702D, is provided in the $\delta T$ column.
@@ Appendix B Log of Trigger Times
Einstein Probe EP250702aaaThis is the start of the initial detection window for the July 1 detection by EP (EP250702a-arxiv) 2025-07-01 01:40:26 60857.0697 EP250702a-arxiv
@@ Appendix C Temporal Fit Results
Figure 14: Top Left: Results of our temporal fit with the afterglow start time free. The fit includes all XRT and Chandra data. Top Right: Results of our temporal fit with the start time $T_{0}$ fixed to GRB 250702D’s trigger time. Bottom Left: Same as top left, but only including XRT data after 0.55 days. Bottom Right: Same as bottom right, but only including XRT data after 0.55 days. The posterior distributions on $T_{0}$ are also shown in Figure 1.
Figure 15: Results of our temporal fit with a broken powerlaw with the onset time $T_{0}$ as a free parameter.